# Knowledge Graph Based Agent for Complex, Knowledge-Intensive QA in Medicine

### v1

@@ B.3 Implementation Details of Baseline Models
B.3.1 Implementation Details of Baseline Models under Multi-Choice Setting
@@
We evaluate these LLMs across all datasets in two distinct settings: the traditional inference setting and the CoT inference setting, using the Transformers package (v 4.31.1) with default parameters. The prompts used for evaluation are provided in C.1.
@@
B.3.2 Implementation Details of Baseline Models under Open-ended Reasoning Setting
To test baseline models under the open-ended reasoning setting, we add a new module for both LLM-based reasoning models and RAG-based models. In terms of KG-based models, they could not be deployed under the open-ended reasoning setting due to their model architectures, as shown in C.2.
For each model in the other two groups, we first get a response to the input descriptive query and then adopt the same LLM as their original design to match the response to the content of the correct answer without considering the input question.
B.4 Implementation Details of Sensitivity Analysis
B.4.1 Effect of Answer Order/Index
@@
B.5 Implementation Details of Adaptability Analysis
@@
Appendix C Prompts
C.1 Prompt Template for Evaluating Baseline Models Under Multi-choice Reasoning Setting
@@
C.2 Prompt Template for Evaluating Baseline Models Under Open Reasoning Setting
@@
C.3 Prompt Template in KGARevion
@@
Given the following question stem, medical terminologies, and options, generate a set of related undirected triplets. Each triplet should consist of a head entity, a relation, and a tail entity. The relations should describe meaningful interactions or associations between the entities, particularly in a medical or biomedical context. Use the query stem and the medical entities contained each option to extract triplets that are relevant to the query and can answer the query correctly. Each triplet should be in the format: (Head Entity, Relationship, Tail Entity). Since the triplets are undirected, the order of Head Entity and Tail Entity does not imply any directional relationship between them. The relationship should be one of the following: [’protein__\__protein’, ’carrier’, ’enzyme’, ’target’, ’transporter’, ’contraindication’, ’indication’, ’off-label use’, ’synergistic interaction’, ’associated with’, ’parentchild’, ’phenotype absent’, ’phenotype present’, ’side effect’, ’interacts with’, ’linked to’, ’expression present’, ’expression absent’]. Ensure that each entity in the triplet is specific and concise, such as diseases, proteins, conditions, symptoms, drugs, treatments, anatomical parts, or other relevant medical entities.
@@
Appendix D Tables
D.1 Description Template
@@ D       The description dictionary for relationship r
      $d_{p}$       The dimension of token embeddings in LLM

### v2

@@ C.3 Implementation Details of Baseline Models
C.3.1 Implementation Details of Baseline Models under Multi-Choice Setting
@@
We evaluate these LLMs across all datasets in two distinct settings: the traditional inference setting and the CoT inference setting, using the Transformers package (v 4.31.1) with default parameters. The prompts used for evaluation are provided in E.1.
@@
C.3.2 Implementation Details of Baseline Models under Open-ended Reasoning Setting
To test baseline models under the open-ended reasoning setting, we add a new module for both LLM-based reasoning models and RAG-based models. In terms of KG-based models, they could not be deployed under the open-ended reasoning setting due to their model architectures, as shown in E.2.
For each model in the other two groups, we first get a response to the input descriptive question and then adopt the same LLM as their original design to match the response to the content of the correct answer without considering the input question.
C.4 Implementation Details of Sensitivity Analysis
C.4.1 Effect of Answer Order/Index
@@
C.5 Implementation Details of Adaptability Analysis
@@
Appendix D KGARevion
D.1 Implementation Details in Revise Action
The Revise action is carried out by providing an instruction to an unfine-tuned LLM, such as LLaMA3.1-8B, instead of retraining the model. The instruction directs the LLM to correct the ”False” triplet identified by the Review action by modifying either the head or tail entity.
  • Input of False Triplets The Revise action begins by receiving False triplets, such as (HSPA1A, interactions, DHDDS), which were flagged as False by the Review action.
  • Instruction to the LLM The system then provides an instruction to the LLM, telling it that the given triplet is incorrect. The instruction asks the LLM to revise the triplet, replacing either the head or tail entity, and to generate a corrected triplet that is relevant to the input question. For example, the instruction could be: ”The following triplet is incorrect: (HSPA1A, interactions, DHDDS). Please revise it to a correct triplet related to the input query.” The details of Instruction is available in E.3
[…]
  • Review of Revised Triplet Once the revised triplet is generated, KGARevion sends it to the Review action for validation. The Review action checks the correctness of the newly generated triplet.
  • Iterative Revision Process If the revised triplet is identified as True by the Review action, it is passed on to the Answer action for final output. If the revised triplet is still identified as False, it is sent back to the Revise action for further modification. This process repeats until either a True triplet is identified or the maximum number of revision rounds is reached.
D.2 Implementation Details in Answer Action
The Answer action is performed by prompting an unfine-tuned LLM, such as LLaMA3.1-8B, instead of retraining the model. This approach mirrors the Revise action, allowing the system to leverage pre-trained models without the need for additional training.
  • Input of Verified Triplets The input to the Answer action consists of all triplets that have been verified as True by the Review actions.
  • Instruction to the LLM Once the verified triplets are received, KGARevion sends an instruction to the LLM. This instruction directs the LLM to generate the final answer based on the provided triplets.
  • Final Answer Generation The LLM receives all True triplets and outputs the final answer based on those verified True triplets.
Appendix E Prompts
E.1 Prompt Template for Evaluating Baseline Models Under Multi-choice Reasoning Setting
@@
E.2 Prompt Template for Evaluating Baseline Models Under Open Reasoning Setting
@@
E.3 Prompt Template in KGARevion
@@
Given the following question stem, medical terminologies, and options, generate a set of related undirected triplets. Each triplet should consist of a head entity, a relation, and a tail entity. The relations should describe meaningful interactions or associations between the entities, particularly in a medical or biomedical context. Use the question stem and the medical entities contained each option to extract triplets that are relevant to the query and can answer the query correctly. Each triplet should be in the format: (Head Entity, Relationship, Tail Entity). Since the triplets are undirected, the order of Head Entity and Tail Entity does not imply any directional relationship between them. The relationship should be one of the following: [’protein__\__protein’, ’carrier’, ’enzyme’, ’target’, ’transporter’, ’contraindication’, ’indication’, ’off-label use’, ’synergistic interaction’, ’associated with’, ’parentchild’, ’phenotype absent’, ’phenotype present’, ’side effect’, ’interacts with’, ’linked to’, ’expression present’, ’expression absent’]. Ensure that each entity in the triplet is specific and concise, such as diseases, proteins, conditions, symptoms, drugs, treatments, anatomical parts, or other relevant medical entities.
@@
Appendix F Tables
F.1 Description Template
@@ D       The description dictionary for relationship r
      $d_{L}$       The dimension of token embeddings in LLM